# Identifiable Causal Inference with Noisy Treatment and No Side Information

### v2

@@ 2.2.1 Model training
(a) Causal graph assumed by CE-oracle, which has access to the true treatment $X^{*}$, unlike the other methods.
@@
(b) Causal graph (incorrectly) assumed by CE-simple.
@@ 2.3.1 Identifiability of causal estimation with a noisy treatment
The measurement error model in Section 1 is model identifiable if 1) for every $z$, $\mu_{Y}(z,x^{*})$ is continuously differentiable everywhere as a function of $x^{*}$, 2) for every $z$, the set $\chi=\{x^{*}:\frac{\partial}{\partial x^{*}}\mu_{Y}(z,x^{*})=0\}$ has at most a finite number of elements, and 3) there exists $z$ for which $\mu_{Y}(z,x^{*})$ is not linear in $x^{*}$ (i.e. of the form $\mu_{Y}(x^{*})=ax^{*}+b$).
@@ 3.1.1 Synthetic datasets from Gaussian processes
For the synthetic experiment, we generate datasets whose underlying distribution follows the CEME model with all variables being scalars. Gaussian processes (GPs) are used, see Rasmussen & Williams (2006) for an introduction. In brief, however, a random function $f(t)$ for $t\in{\mathbb{R}}^{n}$ is a Gaussian process if $f(t_{1}),...,f(t_{k})$ is jointly Gaussian for any $k\in{\mathbb{N}}$ and any $t_{1},...,t_{k}\in{\mathbb{R}}^{n}$. They are denoted by $\textrm{GP}(\mu,K)$ and characterized by a mean function $\mu(t_{i})$ and kernel function $K(t_{i},t_{j})$, by which one obtains the mean vector and covariance matrix for the jointly Gaussian $f(t_{1}),...,f(t_{k})$ (for any $k\in{\mathbb{N}}$ and any $t_{1},...,t_{k}\in{\mathbb{R}}^{n}$). In this study, the mean functions $\mu(t)$ are constant. The data generation proceeds according to Algorithm 1.
@@
2:Sample $\mu_{X^{*}}$ approximately from $\textrm{GP}(0,K)$ and $g^{-1}\circ\sigma_{X^{*}}$ approximately from $\textrm{GP}(1,K)$. Here $\circ$ denotes function composition and $g(z)=\log(1+\exp(z))$ is used to ensure the positivity of $\sigma_{X^{*}}$.
@@
4:Sample $\mu_{Y}$ approximately from $\textrm{GP}((0,0),K)$.
@@
For the GPs, we use the squared exponential kernel $K(u_{1},u_{2})=\alpha\exp\left(-\frac{|u_{1}-u_{2}|^{2}}{2l^{2}}\right)$ with $\alpha=1$ and lengthscale $l=2$ and where |.||\ .\ || . | denotes the Euclidean norm. Note that for $\mu_{X^{*}}(z)$ and $\sigma_{X^{*}}(z)$ the corresponding GPs are over ${\mathbb{R}}$ while for $\mu_{Y}(z,x^{*})$ the GP is over ${\mathbb{R}}^{2}$. The approximations in steps 2 and 4 are done to avoid excessive computational cost that otherwise results from kernel matrices growing proportionally to the square of the number of data points. This is described, along with other experiment details, in Appendix B. Three datasets generated in this manner are shown in Figure 4.
The constant $L$ is varied to obtain datasets with different levels of noise, which are either $L=0.1$ (small noise), $L=0.2$ (medium noise), or $L=0.4$ (large noise). The different training dataset sizes used are 1000, 4000, and 16000 data points. The test data (used for evaluating the models) consists of 20000 data points. For each combination of noise level and training dataset size, we generate 200 different datasets (with different underlying distributions) using Algorithm 1. For all of these, we fit each model (CEME, CEME+, CE-simple and CE-oracle) 6 times, and for the results pick the run with the best score on a separate validation dataset (distinct from the test dataset). All models use the same neural network architecture to facilitate a fair comparison, i.e. three fully connected hidden layers with 20 nodes each and the ELU activation function.
Figure 5: Error in ${\mathbb{E}}[y|z,do(x^{*})]$ estimation in the synthetic experiment. In addition to data for CE-simple, missing from the figure are some outliers for CEME and CEME+ for noise level 40% and training dataset size 1000.
@@ 3.1.2 Results
Figure 6: Error in estimation of $\Delta Y$ standard deviation in the synthetic experiment. In addition to data for CE-simple, missing from the figure are 4 outliers for CEME, 1000 training data and 10% noise as well as 2 outliers for CEME, 4000 training data and 10% noise.
@@
Overall the CEME algorithms offer significant benefit over not taking measurement error into account at all (algorithm CE-simple), both in terms of estimating individual causal effect (Figures 5 and 6), and average causal effect (Figure 7). In addition, the results suggest convergence of estimates when increasing dataset size, as expected from the theoretical identifiability analysis. On the other hand, even with knowledge of the true standard deviation (SD) of $\Delta X$, CEME+ does not achieve performance comparable to that of CE-oracle. This is to be expected as CE-oracle sees accurate values of $X^{*}$ for individual data points, which cannot be identified even in principle from information available to CEME and CEME+.
Figure 7: Error in estimation of $p(y|do(x^{*}))$ in the synthetic experiment. Besides data for CE-simple, missing from the figure are 3 outliers for CEME and 1 for CEME+, both for 40% noise and 1000 training data.
Figure 8: Error in estimation of $\Delta X$ standard deviation by CEME in the synthetic experiment.
@@ 3.2 Experiment with education-wage data
To evaluate our methods, we need to know the ground truth causal effect, and this is why we first train a neural network to predict the outcome, and then modify the dataset by replacing the outcome values with the neural network predictions to which Gaussian noise is added. The noisy treatment $X$ is obtained by adding Gaussian noise to $X^{*}$. Six separate datasets are created, each corresponding to a different level of SD of $\Delta X$. The levels are proportional to the SD of $X^{*}$, and are 0%, 20%, 40%, 60%, 80% and 100%. The choice of education years as treatment was in part motivated by the need to know the true treatment accurately to obtain the ground truth causal effect. The CEME/CEME+ models are misspecified for this dataset because the true treatment, the number of education years, is ordinal, instead of conditionally Gaussian, as the models assume. This presents an opportunity to evaluate the sensitivity of the CEME/CEME+ algorithms to model misspecification.
@@ 3.2.1 Results
Figure 9: Error in causal effect estimation in the education-wage data experiment.
@@
We notice that while CEME and CEME+ still offer a clear improvement over not accounting for measurement at all (CE-simple), they seem to be further below in performance from CE-oracle (which acts as a benchmark for optimal, or even beyond-optimal performance for the CEME/CEME+ algorithms). A potential reason for this is that the CEME models, which model the true number of education years as a continuous latent variable, are misspecified unlike CE-oracle and CE-simple, which only condition on the number of education years but do not model it (though CE-simple assumes that there is no measurement error). Moreover, CEME and CEME+ have more parameters than CE-oracle and CE-simple; therefore they could be hurt more by the limited dataset size and high-dimensional covariate. The larger difference between CEME and CEME+ than in the synthetic experiment might be because CEME+ avoids some of the detriment from model misspecification by having access to the true SD of $\Delta X$.
@@
Figure 10: Error in the estimation of treatment and outcome noises in the education-wage data experiment.
@@ 4 Conclusion
In this study, we provided a model for causal effect estimation in the presence of treatment noise, with complex nonlinear dependencies, and with no available side information. We confirmed the model’s identifiability theoretically and evaluated it experimentally, comparing it to a baseline that does not account for measurement error at all (CE-simple) and to an upper bound in performance in the form of CE-oracle. A notable advantage of the model is its flexibility: It offered good performance on a diverse set of synthetic datasets and was useful for correcting for measurement error even on a real-world dataset for which it was clearly misspecified. Our approach is also flexible in a second way: after using amortized variational inference to infer the SCM and the posterior of hidden variables, any interventional or counterfactual distributions may be computed.
@@ B.1 Synthetic experiment
Learning rate for CE-oracle and CE-simple 0.001
@@ Appendix D Proof of Proposition 1
First, consider the measurement error model in Section 1 conditioned on a specific value $Z=z$, effectively removing $Z$ as a variable. We show that this model, called the restricted model, as opposed to the original full model, satisfies the assumptions in Theorem 1, which thus determines when the restricted model is identifiable. First, the restricted model satisfies Equations (14) and (15). Assumption 1 follows directly from the definition of the model, as $|\Delta X|$ and $|\Delta Y|$ follow the half-normal distribution, which has the known finite expectation $\sigma\sqrt{2}/\sqrt{\pi}$. Assumption 2 is satisfied because $\Delta X$ and $\Delta Y$ are Gaussian so their characteristic functions have the known form $\exp(i\mu t-\sigma^{2}t^{2}/2)$ and thus do not vanish for any $t\in{\mathbb{R}}$ ($t$ is denoted by $\xi$ or $\gamma$ in Assumption 2).

### v3

@@ 2.2.1 Model training
(a) Causal graph assumed by Oracle, which has access to the true treatment $X^{*}$, unlike the other methods.
@@
(b) Causal graph (incorrectly) assumed by the method Naive.
@@ 2.3.1 Identifiability of causal estimation with a noisy treatment
The measurement error model defined in Equations (1)–(5) is model identifiable if 1) for every $z$, $\mu_{Y}(z,x^{*})$ is continuously differentiable everywhere as a function of $x^{*}$, 2) for every $z$, the set $\chi=\{x^{*}:\frac{\partial}{\partial x^{*}}\mu_{Y}(z,x^{*})=0\}$ has at most a finite number of elements, and 3) there exists $z$ for which $\mu_{Y}(z,x^{*})$ is not linear in $x^{*}$ (i.e. of the form $\mu_{Y}(x^{*})=ax^{*}+b$).
@@ 3.1.1 Synthetic datasets from Gaussian processes
For the synthetic experiment, we generate datasets whose underlying distribution follows the CEME model with all variables being scalars. Gaussian processes (GPs) are used, see Rasmussen & Williams (2006) for an introduction. In brief, a random function $f(t)$ for $t\in{\mathbb{R}}^{n}$ is a Gaussian process if $f(t_{1}),...,f(t_{k})$ is jointly Gaussian for any $k\in{\mathbb{N}}$ and any $t_{1},...,t_{k}\in{\mathbb{R}}^{n}$. They are denoted by $\textrm{GP}(\mu,K)$ and characterized by a mean function $\mu(t_{i})$ and kernel function $K(t_{i},t_{j})$, by which one obtains the mean vector and covariance matrix for the jointly Gaussian $f(t_{1}),...,f(t_{k})$ (for any $k\in{\mathbb{N}}$ and any $t_{1},...,t_{k}\in{\mathbb{R}}^{n}$). In this study, the mean functions $\mu(t)$ are constant. The data generation proceeds according to Algorithm 1.
@@
2:Sample $\mu_{X^{*}}$ from $\textrm{GP}(0,K)$ and $g^{-1}\circ\sigma_{X^{*}}$ from $\textrm{GP}(1,K)$. Here $\circ$ denotes function composition and $g(z)=\log(1+\exp(z))$ is used to ensure the positivity of $\sigma_{X^{*}}$.
@@
4:Sample $\mu_{Y}$ from $\textrm{GP}((0,0),K)$.
@@
For the GPs, we use the squared exponential kernel $K(u_{1},u_{2})=\alpha\exp\left(-\frac{|u_{1}-u_{2}|^{2}}{2l^{2}}\right)$ with $\alpha=1$ and lengthscale $l=2$ and where |.||\ .\ || . | denotes the Euclidean norm. Note that for $\mu_{X^{*}}(z)$ and $\sigma_{X^{*}}(z)$ the corresponding GPs are over ${\mathbb{R}}$ while for $\mu_{Y}(z,x^{*})$ the GP is over ${\mathbb{R}}^{2}$. To avoid excessive computational cost, the functions $\mu_{X^{*}}(z)$, $\sigma_{X^{*}}(z)$ and $\mu_{Y}(z,x^{*})$ are only approximations of true samples from the GPs (steps 2 and 4 in Algorithm 1), as described in Appendix B along with other experiment details. Three datasets generated in this manner are shown in Figure 4.
The constant $L$ is varied to obtain datasets with different levels of noise, which are either $L=0.1$ (small noise), $L=0.2$ (medium noise), or $L=0.4$ (large noise). The different training dataset sizes used are 1000, 4000, and 16000 data points. The test data (used for evaluating the models) consist of 20000 data points. For each combination of noise level and training dataset size, we generate 200 different datasets (with different underlying distributions) using Algorithm 1. For all of these, we fit each model (CEME, CEME+, Naive and Oracle) 6 times, and for the results pick the run with the best score on a separate validation dataset (distinct from the test dataset). All models use the same neural network architecture to facilitate a fair comparison, i.e. three fully connected hidden layers with 20 nodes each and the ELU activation function.
Figure 5: Error in ${\mathbb{E}}[y|z,do(x^{*})]$ estimation in the synthetic experiment. In addition to data for Naive, missing from the figure are some outliers for CEME and CEME+ for noise level 40% and training dataset size 1000. Key observations are that 1) the proposed CEME/CEME+ methods offer clear benefit over Naive that does not account for measurement error, 2) CEME and CEME+ seem to converge with increasing training set size and compare relatively well with the loose upper bound Oracle, and 3) CEME handles unknown measurement error variance well, since CEME+ that knows it, performs only slightly better.
@@ 3.1.2 Results
Figure 6: Error in estimation of $\Delta Y$ standard deviation in the synthetic experiment. In addition to data for Naive, missing from the figure are 4 outliers for CEME, 1000 training data and 10% noise as well as 2 outliers for CEME, 4000 training data and 10% noise. Key observations are that 1) the estimates seem to converge for the proposed CEME/CEME+ but not for the baseline Naive, and 2) all the methods tend to overestimate, rather than underestimate, the standard deviation of $\Delta Y$.
@@
Overall the CEME algorithms offer significant benefit over not taking measurement error into account at all (algorithm Naive), both in terms of estimating individual causal effect (Figures 5 and 6), and average causal effect (Figure 7). In addition, the results suggest convergence of estimates when increasing dataset size, as expected from the theoretical identifiability analysis. On the other hand, even with knowledge of the true standard deviation (SD) of $\Delta X$, CEME+ does not achieve performance comparable to that of Oracle. This is to be expected as Oracle sees accurate values of $X^{*}$ for individual data points, which cannot be identified even in principle from information available to CEME and CEME+.
Figure 7: Error in estimation of $p(y|do(x^{*}))$ in the synthetic experiment. Besides data for Naive, missing from the figure are 3 outliers for CEME and 1 for CEME+, both for 40% noise and 1000 training data. This figure tells a similar story to Figure 5, except that the differences between the methods are more pronounced.
Figure 8: Error in estimation of $\Delta X$ standard deviation by CEME in the synthetic experiment. The estimates seem to converge with increasing training set size.
@@ 3.2 Experiment with education-wage data
To evaluate our methods, we need to know the ground truth values of the parameters that we estimate. Having access to real observed data alone does not permit this, which is why we augment the real data with synthetic variables that mimic the real ones. The known data generating processes of these synthetic variables enable us to compute the ground truth. To this end, we first train a neural network to predict the outcome, and then modify the dataset by replacing the outcome values with the neural network predictions to which Gaussian noise is added. The noisy treatment $X$ is obtained by adding Gaussian noise to $X^{*}$. Six separate datasets are created, each corresponding to a different level of SD of $\Delta X$. The levels are proportional to the SD of $X^{*}$, and are 0%, 20%, 40%, 60%, 80% and 100%. The choice of education years as treatment was in part motivated by the need to know the true treatment accurately to obtain the ground truth causal effect. The CEME/CEME+ models are misspecified for this dataset because the true treatment, the number of education years, is ordinal, instead of conditionally Gaussian, as the models assume. This presents an opportunity to evaluate the sensitivity of the CEME/CEME+ algorithms to model misspecification.
@@ 3.2.1 Results
Figure 9: Error in causal effect estimation in the education-wage data experiment. The proposed CEME/CEME+ methods offer improvement compared to the baseline Naive, but not as much as in the synthetic experiment. A likely reason for this is that CEME/CEME+ incorrectly assume a continuous $X^{*}$, unlike Oracle and Naive. Moreover, the difference between CEME and CEME+ is larger than in the synthetic case.
@@
We notice that while CEME and CEME+ still offer a clear improvement over not accounting for measurement at all (Naive), they seem to be further below in performance from Oracle (which acts as a benchmark for optimal, or even beyond-optimal performance for the CEME/CEME+ algorithms). A potential reason for this is that the CEME models, which model the true number of education years as a continuous latent variable, are misspecified unlike Oracle and Naive, which only condition on the number of education years but do not model it (though Naive assumes that there is no measurement error). Moreover, CEME and CEME+ have more parameters than Oracle and Naive; therefore they could be hurt more by the limited dataset size and high-dimensional covariate. The larger difference between CEME and CEME+ than in the synthetic experiment might be because CEME+ avoids some of the detriment from model misspecification by having access to the true standard deviation of $\Delta X$.
@@
Figure 10: Error in the estimation of treatment and outcome noises in the education-wage data experiment. On the left, only CEME is depicted because it is the only method that estimates $\Delta X$ standard deviation. These figures help demonstrate that the proposed CEME/CEME+ offer clear benefit over the baseline Naive even when they are misspecified.
@@ 4 Conclusion
In this study, we provided a model for causal effect estimation in the presence of treatment noise, with complex nonlinear dependencies, and with no side information or knowledge of the measurement error variance. We confirmed the model’s identifiability theoretically and evaluated it experimentally, comparing it to a baseline that does not account for measurement error at all (Naive) and to an upper bound in performance in the form of Oracle. A notable advantage of the model is its flexibility: It offered good performance on a diverse set of synthetic datasets and was useful for correcting for measurement error even on a real-world dataset for which it was clearly misspecified. Our approach is also flexible in a second way: after using amortized variational inference to infer the SCM and the posterior of hidden variables, any interventional or counterfactual distributions may be computed.
@@
Broader Impact Statement
Our method could be used for estimating treatment effects in applications where incorrect inferences might in the worst case lead to severe adverse outcomes, e.g. in healthcare. For this reason it is important to consider the validity of the assumptions of our method in any particular use case as well as to test any implementations carefully before practical deployment.
Author Contributions
All authors have accepted responsibility for the entire content of this manuscript and approved to its submission. AP had the main responsibility on all aspects of the work. PM supervised the work and provided the initial research idea. The authors wrote the paper together.
Acknowledgments
This work was supported by the Academy of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI, and grants 336033, 352986, 358246) and EU (H2020 grant 101016775 and NextGenerationEU).
@@ B.1 Synthetic experiment
Number of hidden layers in each network (fully connected) 3
Width of each hidden layer 20
Activation function ELU
@@
Learning rate for Oracle and Naive 0.001
@@ B.2 Experiment with education-wage data
Number of hidden layers in each network (fully connected) 3
Width of each hidden layer 26
Activation function ELU
@@ Appendix D Proof of Proposition 1
First, consider the measurement error model defined in Equations (1)–(5) conditioned on a specific value $Z=z$, effectively removing $Z$ as a variable. We show that this model, called the restricted model, as opposed to the original full model, satisfies the assumptions in Theorem 1, which thus determines when the restricted model is identifiable. First, the restricted model satisfies Equations (13) and (14). Assumption 1 follows directly from the definition of the model, as $|\Delta X|$ and $|\Delta Y|$ follow the half-normal distribution, which has the known finite expectation $\sigma\sqrt{2}/\sqrt{\pi}$. Assumption 2 is satisfied because $\Delta X$ and $\Delta Y$ are Gaussian so their characteristic functions have the known form $\exp(i\mu t-\sigma^{2}t^{2}/2)$ and thus do not vanish for any $t\in{\mathbb{R}}$ ($t$ is denoted by $\xi$ or $\gamma$ in Assumption 2).